### v1

# Calibrating dimension reduction hyperparameters in the presence of noise

Abstract
The goal of dimension reduction tools is to construct a low-dimensional representation of high-dimensional data. These tools are employed for a variety of reasons such as noise reduction, visualization, and to lower computational costs. However, there is a fundamental issue that is highly discussed in other modeling problems, but almost entirely ignored in the dimension reduction literature: overfitting. If we interpret data as a combination of signal and noise, prior works judge dimension reduction techniques on their ability to capture the entirety of the data, i.e. both the signal and the noise. In the context of other modeling problems, techniques such as feature-selection, cross-validation, and regularization are employed to combat overfitting, but no such precautions are taken when performing dimension reduction. In this paper, we present a framework that models dimension reduction problems in the presence of noise and use this framework to explore the role perplexity and number of neighbors play in overfitting data when applying t-SNE and UMAP. We also present a workflow others may use to calibrate perplexity or number of neighbors in the presence of noise.
1 Introduction
Since the introduction of t-SNE, perplexity calibration has proven to be a difficult task, so much so that perplexity-free versions of t-SNE have been proposed (Crecchi et al., 2020). It is also an extremely important task, since t-SNE is known to produce unfaithful results when mishandled (Huang et al., 2022). The original authors suggested values between 5 and 50 (van der Maaten & Hinton, 2008), while recent works have suggested perplexities as large as one percent of the sample size (Kobak & Berens, 2019). (Cao & Wang, 2017) studied the inverse relationship between perplexity and Kullback-Leibler divergence to design an automatic calibration process that “generally agrees with experts’ consensus.” Manual tuning of perplexity requires a deep understanding of the t-SNE algorithm and general knowledge of the data’s structure. When the data’s structure is available, we can visualize the results and choose the perplexity that best captures the hypothesized structure. In supervised problems, for example, we look for low-dimensional representations that cluster according to the class labels. For unsupervised problems, however, the structure is often unknown, so we cannot visually assess each representation. In these cases, we must resort to quantitative measures of performance to understand how the well the low-dimensional representation represents the high-dimensional data. While this strategy is heavily discussed in the machine learning literature, prior works disregard the possibility of overfitting when measuring performance.
In this paper, we present a framework for studying dimension reduction methods in the presence of noise (Section 3). We then use this framework to calibrate t-SNE and UMAP hyperparameters in both simulated and practical examples to illustrate how the disregard of noise leads to miscalibration (Section 4). We also discuss how other researchers may use this framework in their own work (Section 5).
2 Background
2.1 t-SNE
t-distributed stochastic neighbor embedding (t-SNE) is a nonlinear dimension reduction method primarily used for visualizing high-dimensional data. The t-SNE algorithm captures the topological structure of high-dimensional data by calculating directional similarities via a Gaussian kernel. The similarity of point $x_{j}$ to point $x_{i}$ is defined by
$p_{j|i}=\frac{\exp(-||x_{i}-x_{j}||^{2}/2\sigma_{i}^{2})}{\sum_{k\neq i}\exp(-||x_{i}-x_{k}||^{2}/2\sigma_{i}^{2})}.$
Thus for each point $x_{i}$, we have a a probability distribution $P_{i}$ that quantifies the similarity of every other point to $x_{i}$. The variance of the Gaussian kernel $\sigma_{i}^{2}$ is chosen so that the perplexity of the probability distribution $P_{i}$, in the information theory sense, is equal to a pre-specified value also named perplexity,
$\textrm{perplexity}=2^{-\sum_{j\neq i}p_{j|i}\log_{2}p_{j|i}.}$
Intuitively, perplexity controls how large a neighborhood to consider around each point when approximating the topological structure of the data (van der Maaten and Hinton, 2008). As such, it implicitly balances attention to local and global aspects of the data with high values of perplexity placing more emphasis on global aspects. For the sake of computational convenience, t-SNE assumes the directional similarities are symmetric by defining
$p_{ij}=\frac{p_{i|j}+p_{j|i}}{2n}.$
The $p_{ij}$ define a probability distribution $P$ on the set of pairs $(i,j)$ that represents the topological structure of the data.
The goal is to then find an arrangement of low-dimensional points $y_{1},\ldots,y_{n}$ whose similarities $q_{ij}$ best match the $p_{ij}$ in terms of Kullback-Leibler divergence,
$D_{KL}(P||Q)=\sum_{i,j}p_{ij}\log\frac{p_{ij}}{q_{ij}}.$
The low-dimensional similarities $q_{ij}$ are defined using the t-distribution with one degree of freedom,
$q_{ij}=\frac{(1+||y_{i}-y_{j}||^{2})^{-1}}{\sum_{k\neq l}(1+||y_{k}-y_{l}||^{2})^{-1}}.$
The main downsides of t-SNE are its inherit randomness and sensitivity to hyperparameter calibration. The minimization of the KL divergence is done using gradient descent methods with incorporated randomness to avoid stagnating at local minima. As a result, the output differs between runs of the algorithm. Hence, the traditional t-SNE workflow often includes running the algorithm multiple times at various perplexities before choosing the best representation.
2.2 UMAP
Uniform Manifold Approximation and Projection (UMAP) is another nonlinear dimension reduction method that has been rising in popularity. Like t-SNE, UMAP is a powerful tool for visualizing high-dimensional data that requires user calibration. The architecture of the UMAP algorithm is similar to that of t-SNE’s — high-dimensional similarities are computed and the outputted representation is the set of low-dimensional points whose low-dimensional similarities best match the high-dimensional similarities. The cost function and the formulas for high/low-dimensional similarities differ from t-SNE. See (McInnes et al., 2020) for details.
UMAP has a hyperparameter called n_neighbors that is analogous to t-SNE’s perplexity. It determines how large a neighborhood to use around each point when calculating high-dimensional similarities. The original authors make no recommendation for optimal values of n_neighbors, but their implementation defaults to n_neighbors = 15 (McInnes et al., 2020).
3 Methods
3.1 Dimension Reduction Framework
Prior works quantitatively measure how well low-dimensional representations match the high-dimensional data. However, if we view the original data as a composition of signal and noise, we must not reward capturing the noise. This would lead to overfitting. Therefore, we should be comparing the low-dimensional representation against the signal underlying our data, rather than the entirety of the data.
Suppose the underlying signal of our data is captured by an $r$-dimensional matrix $Y\in\mathbb{R}^{n\times r}$. In the context of dimension reduction, the underlying signal is often lower dimension than the original data. Let $p\geq r$ be the dimension of the original data set, and let $\textrm{Emb}:\mathbb{R}^{r}\to\mathbb{R}^{p}$ be the function that embeds the signal in data space. Define $Z=\textrm{Emb}(Y)$ to be the signal embedded in data space. We then assume the presence of random error. The original data can then be modeled by
$Z+\epsilon\textrm{ for }\epsilon\sim N_{p}(0,\Sigma).$
The dimension reduction method $\varphi$ is applied to $Z+\epsilon$ to get a low-dimensional representation $X\in\mathbb{R}^{n\times q}$. See Figure 1.
{tikzcd}
Figure 1: Dimension reduction framework
3.2 Reconstruction Error Functions
The remaining piece is a procedure for measuring dimension reduction performance. Suppose we have a reconstruction error function $f(D_{1},D_{2})$ that quantifies how well the data set $D_{2}$ represents the data set $D_{1}$. Prior works like (Chari & Pachter, 2023), (Espadoto et al., 2021), (Huang et al., 2022), and (Kodak & Berens, 2019) use various reconstruction error functions to quantify performance; only, they study $f(Z+\epsilon,X)$ to measure how well the constructed representation $X$ represents the original data $Z+\epsilon$. We argue it is more appropriate to compare $X$ against the underlying signal $Y$ by examining $f(Y,X)$.
Prior works in dimension reduction have suggested various quantitative metrics for measuring dimension reduction performance. In line with recent discussions of perplexity [(Kodak & Berens, 2019) and (Chari & Pachter, 2023)], we focus on two different metrics — one that measures local performance and one that measures global performance.
For local performance, we use a nearest-neighbor type metric called trustworthiness (Venna and Kaski, 2006). Let $n$ be the sample size and $r(i,j)$ the rank of point $j$ point among the $k$ nearest neighbors of point $i$ in high dimension. Let $U_{k}(i)$ denote the set of points among the $k$ nearest neighbors of point $i$ in low dimension, but not high dimension. Then
$f_{trust}(D_{1},D_{2})=1-\frac{2}{nk(2n-3k-1)}\sum_{i=1}^{n}\sum_{j\in U_{k}(i)}\left[r(i,j)-k\right].$
For each point, we are measuring the degree of intrusion into its $k$-neighborhood during the dimension reduction process. The quantity is then re-scaled, so that trustworthiness falls between 0 and 1 with higher values favorable. Trustworthiness is preferable to simply measuring the proportion of neighbors preserved because it’s more robust to the choice of $k$. For very large values of $n$, we can get an estimate by only checking a random subsample of points $i_{1},\ldots,i_{m}$. In this case,
$f_{trust}(D_{1},D_{2})\approx 1-\frac{2}{mk(2n-3k-1)}\sum_{l=1}^{m}\sum_{j\in U_{k}(i_{l})}\left[r(i_{l},j)-k\right].$
For global performance, we use Shepard goodness (Espadoto et al., 2021). Shepard goodness is the Kendall correlation (a rank-based correlation) between high and low-dimensional inter-point distances,
$f_{\textrm{Shep}}(D_{1},D_{2})=\sigma_{\textrm{Kendall}}(||z_{i}-z_{j}||^{2},||\varphi(z_{i})-\varphi(z_{j})||^{2}).$
Again for very large values of $n$, we can get an approximation by calculating the correlation between inter-point distances of a random subsample.
3.3 Using this framework
When using our framework to simulate examples, three components must be specified: $Z+\epsilon$, $Y$, and $\textrm{Emb}()$. These elements describe the original data, the underlying signal, and the embedding of the signal in data space, respectively. When simulating examples, it’s natural to start with the underlying signal $Y$ then construct $Z+\epsilon$ by attaching extra dimensions and adding Gaussian noise. The $\textrm{Emb}()$ function is then given by
$\textrm{Emb}(y)=(y,0,\ldots,0),$
so that
$Z+\epsilon=\begin{bmatrix}Y&|&0\end{bmatrix}+\epsilon.$
Figure 3a: Shepard Goodness vs. Trustworthiness (Links)
Figure 3b: Trustworthiness vs. Perplexity (Links)
Practical examples are more tricky because we do not have the luxury of first defining $Y$. Instead, we are given the data $Z+\epsilon$ from which we must extract $Y$, or at least our best estimate. This process is left to the researcher and should be based on a priori knowledge of the data. If there is no specific signal of interest, a more general approach can be taken. We used a PCA projection of the data to represent the signal,
$Y=\textrm{PCA}_{r}(Z+\epsilon),$
where $r$ is the dimension of the projection. For a reasonable $r$, we would expect the first $r$ principal components to contain most of the signal, while excluding most of the noise. Another advantage to using PCA is it gives rise to a natural $\textrm{Emb}()$ function — the PCA inverse transform. If $Y$ is centered, then we may define
$Z=\textrm{invPCA}_{r}(Y)=(Z+\epsilon)V_{r}V_{r}^{T},$
where $V_{r}\in\mathbb{R}^{p\times r}$ contains the first $r$ eigenvectors of $(Z+\epsilon)^{T}(Z+\epsilon)$ as column vectors.
4 Results
4.1 Simulated Examples
We first looked at simulated examples with explicitly defined signal structures: inter-linked circles, the trefoil knot, and the mammoth dataset (Figure 2).
Figure 2: Simulated Examples; links and trefoil from (Wattenberg et al., 2016), mammoth from (Wang et al., 2021)
Figure 4a: Shepard Goodness vs. Trustworthiness (CyTOF)
Figure 4b: Trustworthiness vs. Perplexity (CyTOF)
For each example, the signal $Y$ consisted of 500 points in three dimensions. The dataset $Z+\epsilon$ was constructed from $Y$ by adding seven superfluous dimensions and isotropic Gaussian noise to all 10 dimensions. We then ran t-SNE using the $R$ package Rtsne with perplexities ranging from 10 to 160333Rtsne implements a Barnes-Hut approximation that only allows for a maximum perplexity of $\frac{n-1}{3}$.. For each value of perplexity, we ran the algorithm 20 times to mimic the ordinary t-SNE workflow. If we were to disregard the distinction between signal and noise, a plot of $f_{\textrm{Shep}}(Z+\epsilon,X)$ vs. $f_{\textrm{trust}}(Z+\epsilon,X)$ could be used to calibrate perplexity. To avoid overfitting the noise, a plot of $f_{\textrm{Shep}}(Y,X)$ vs. $f_{\textrm{trust}}(Y,X)$ should be used instead. See Figure 3a for plots of the links example. Both plots depict an increase in global performance as perplexity increases. Both plots also depict an increase in local performance followed by a decrease as perplexity increases. This trend is more evident when we plot trustworthiness vs. perplexity (Figure 3b).
Past a certain perplexity, local performance begins to deteriorate with increasing perplexity, exhibiting the trade-off between global and local performance. This cutoff point, however, varies between the two plots. When comparing against the original data, a perplexity of 30 maximizes trustworthiness, which is consistent with the original authors’ suggestion of 5 to 50 for perplexity (van der Maaten and Hinton, 2008). When comparing against the signal, a perplexity of 110 maximizes trustworthiness. We hypothesize t-SNE tends to overfit the noise when the perplexity is too low. Intuitively, small perplexities are more affected by slight perturbations of the data when only considering small neighborhoods around each point, leading to unstable representations. Conversely, larger perplexities lead to more stable representations that are less affected by noise. The other two simulated examples exhibit a similar trend (See Supplementary Information).
Figure 5: Trustworthiness vs. Perplexity for $r=5$ (CyTOF)
4.2 Real Examples
For a practical example, we looked at a CyTOF dataset containing 239,933 observations in 49 dimensions (Strauss-Albee et al., 2015). To reduce the computational load, a subset of 5,000 observations was sampled. In line with the ordinary t-SNE workflow, a log transformation was followed by a PCA pre-processing step to reduce the number of dimensions to 10. The processed data set to be studied consisted of 5,000 observations in 10 dimensions,
$Z+\epsilon\in\mathbb{R}^{5,000\times 10}.$
$Y$ was extracted by taking the first three principal components,
$Y=\textrm{PCA}_{3}(Z+\epsilon).$
We computed the t-SNE representations for perplexity values ranging from 20 to 1,600. For each perplexity value, 20 different t-SNE representations were computed. Figure 4a contains the plots for $f_{\textrm{Shep}}(Z+\epsilon,X)$ vs. $f_{\textrm{trust}}(Z+\epsilon,X)$ and $f_{\textrm{Shep}}(Y,X)$ vs. $f_{\textrm{trust}}(Y,X)$.
As with the simulated examples, there is a difference in trend when the frame of reference is the original data versus when it’s just the underlying signal. More specifically, the t-SNE representations better locally represent the original data for lower perplexities and better locally represent the underlying signal for larger perplexities (Figure 4b). When compared against the original data, trustworthiness is maximized at a perplexity of 50, which is consistent with (Kobak and Berens, 2019)’s recommendation of setting perplexity to $n/100$. When compared against the underlying signal, trustworthiness is maximized at a much larger perplexity of 900, reinforcing the hypothesis that lower values of perplexity may be overfitting the noise.
If we, instead, decide to be a little more conservative and use the first five principal components to represent the signal, we still see a similar trend (Figure 5). Trustworthiness still increases then decreases with perplexity. In this case, however, trustworthiness peaks at a perplexity of 100, less than before. By including two extra principal components in the signal, we’re assuming the data contains less noise, allowing the model to be more aggressive during the fitting process. Further experimentation confirmed that the trustworthiness-maximizing perplexity decreased as we included more dimensions in the signal.
4.3 UMAP and n_neighbors
If n_neighbors functions similarly to perplexity, we’d expect small values of n_neighbors to overfit the data as well. Our experiments suggest this to be the case, but to a lesser degree. An identical experiment was run using the Python package umap-learn. n_neighbor values ranging from 20 to 1,000 were tested on the same CyTOF data set. An n_neighbors value of 20 maximized trustworthiness when comparing against the original data, but an n_neighbors value of 600 maximized trustworthiness when comparing against the underlying signal (see Supplementary Information for plots). While the difference in local performance for different values of n_neighbors was smaller, the UMAP results exhibited a similar trend. Local performance peaked at a larger value when compared against the underlying signal versus the original data.
5 Application
To apply this framework in practice, one must decide how to extract the signal from the data. The signal should include the features of the data one desires to retain throughout the dimension reduction process. When using a PCA projection to serve as the signal, one could draw a scree plot or employ a component selection algorithm such as parallel analysis to determine the dimension of the signal (Horn, 1965).
With a signal constructed, it remains to compute t-SNE outputs at varying perplexities. It’s recommended that at least a couple outputs are computed for each perplexity to account for t-SNE’s inherit randomness. For each output, one must calculate the trustworthiness and Shepard goodness with respect to the signal. From there, one can choose the representation with the desirable balance of local and global performance. See algorithm 1. Sample code is available at https://github.com/JustinMLin/DR-Framework/.
Algorithm 1 Measuring Performance in the Presence of Noise
1:original data $Z+\epsilon$, perplexities $\{p_{1},\ldots,p_{m}\}$ to test, and neighborhood size $k$
2:$Y\Leftarrow\textrm{PCA}_{r}(Z+\epsilon)$
3:$\textrm{perplexities}\Leftarrow\{p_{1},\ldots,p_{m}\}$
4:for perplexity in perplexities do
5:     loop
6:         $X\_tsne\Leftarrow\textrm{Rtsne}(Z+\epsilon,\textrm{perplexity})$
7:         $trust\Leftarrow\textrm{trustworthiness}(Y,X\_tsne,k)$
8:         $shep\Leftarrow\textrm{Shepard\_goodness}(Y,X\_tsne)$
9:     end loop
10:end for
11:Plot trustworthiness and Shepard goodness values
12:Choose output with desired balance of local and global performance
It is worth noting that computational barriers may arise, especially for very large data sets. To alleviate such issues, trustworthiness and Shepard goodness can be approximated by subsampling before calculation. Furthermore, t-SNE is generally robust to small changes in perplexity (van der Maaten and Hinton, 2008), so checking a handful of different perplexities is sufficient. If computing the t-SNE representations is the limiting factor, the perplexity can be calibrated for a subsample of the data, instead. (Skrodzki et al., 2023) found that embedding a $\rho$-sample, where $\rho\in(0,1]$ is the sampling rate, with perplexity $\textrm{Perp}^{\prime}$ gives a visual impression of embedding the original data with perplexity
$\textrm{{Perp}}=\frac{\textrm{Perp}^{\prime}}{\rho}.$
With these concessions, applying this framework to calibrate perplexity is feasible for data sets of any size.
6 Discussion
We have illustrated the importance of acknowledging noise when performing dimension reduction by studying the roles perplexity and n_neighbors play in overfitting data. When using the original data to calibrate perplexity, our experiments agreed with perplexities previously recommended. When using just the signal, however, our experiments indicated that larger perplexities performed better. Low perplexities lead to overly-flexible t-SNE models that are heavily impacted by the presence of noise, while higher perplexities exhibit better performance due to increased stability. These considerations are especially important when working with heavily noised data, which are especially prevalent in the world of single-cell transcriptomics (Chu et al., 2022).
We have also presented a framework for modeling dimension reduction problems in the presence of noise. This framework can be used to study other hyperparameters and their relationships with noise. In the case when a specific signal structure is desired, this framework can be used to determine which dimension reduction method best preserves the desired structure. Further works should explore alternative methods for extracting the signal in way that preserves the desired structure.
7 Data Availability
All data and example code are freely available at https://github.com/JustinMLin/DR-Framework/.

### v3

@@
 In this paper, we present a framework that models dimension reduction problems in the presence of noise and use this framework to explore the role perplexity and number of neighbors play in overfitting data when applying t-SNE and UMAP.
-We also present a workflow others may use to calibrate perplexity or number of neighbors in the presence of noise.
+More specifically, we show previously recommended values for perplexity and number of neighbors are too small and tend to overfit the noise.
+We also present a workflow others may use to calibrate hyperparameters in the presence of noise.
 1 Introduction
-Since the introduction of t-SNE, perplexity calibration has proven to be a difficult task, so much so that perplexity-free versions of t-SNE have been proposed (Crecchi et al., 2020).
-It is also an extremely important task, since t-SNE is known to produce unfaithful results when mishandled (Huang et al., 2022).
-The original authors suggested values between 5 and 50 (van der Maaten & Hinton, 2008), while recent works have suggested perplexities as large as one percent of the sample size (Kobak & Berens, 2019).
-(Cao & Wang, 2017) studied the inverse relationship between perplexity and Kullback-Leibler divergence to design an automatic calibration process that “generally agrees with experts’ consensus.” Manual tuning of perplexity requires a deep understanding of the t-SNE algorithm and general knowledge of the data’s structure.
+In recent years, non-linear dimension reduction techniques have been growing in popularity due to their usefulness when analyzing high-dimensional data.
+To date, they are the most effective tools for visualizing high-dimensional data.
+Visualization plays a key role in identifying and understanding relationships within the data.
+It exposes patterns, trends, and outliers that can’t be discerned from the raw data.
+The most popular non-linear dimension reduction technique is t-distributed stochastic neighbor embedding (t-SNE, [17]).
+Since the introduction of t-SNE, hyperparameter calibration has proven to be a difficult task.
+The most crucial hyperparameter, known as perplexity, controls how large a neighborhood to consider around a point when determining its location.
+Perplexity calibration is so troublesome, that perplexity-free versions of t-SNE have been proposed [6].
+It is also an extremely important task, since t-SNE is known to produce unfaithful results when mishandled [9].
+The original authors suggested values between 5 and 50 [17], while recent works have suggested perplexities as large as one percent of the sample size [10].
+[3] studied the inverse relationship between perplexity and Kullback-Leibler divergence to design an automatic calibration process that “generally agrees with experts’ consensus.” Manual tuning of perplexity requires a deep understanding of the t-SNE algorithm and general knowledge of the data’s structure.
 When the data’s structure is available, we can visualize the results and choose the perplexity that best captures the hypothesized structure.
@@
 2.1 t-SNE
-t-distributed stochastic neighbor embedding (t-SNE) is a nonlinear dimension reduction method primarily used for visualizing high-dimensional data.
+t-distributed stochastic neighbor embedding (t-SNE, [17]) is a nonlinear dimension reduction method primarily used for visualizing high-dimensional data.
 The t-SNE algorithm captures the topological structure of high-dimensional data by calculating directional similarities via a Gaussian kernel.
@@
 $p_{j|i}=\frac{\exp(-||x_{i}-x_{j}||^{2}/2\sigma_{i}^{2})}{\sum_{k\neq i}\exp(-||x_{i}-x_{k}||^{2}/2\sigma_{i}^{2})}.$
-Thus for each point $x_{i}$, we have a a probability distribution $P_{i}$ that quantifies the similarity of every other point to $x_{i}$.
-The variance of the Gaussian kernel $\sigma_{i}^{2}$ is chosen so that the perplexity of the probability distribution $P_{i}$, in the information theory sense, is equal to a pre-specified value also named perplexity,
+Thus for each point $x_{i}$, we have a probability distribution $P_{i}$ that quantifies the similarity of every other point to $x_{i}$.
+The scale of the Gaussian kernel, $\sigma_{i}$, is chosen so that the perplexity of the probability distribution $P_{i}$, in the information theory sense, is equal to a pre-specified value also named perplexity,
 $\textrm{perplexity}=2^{-\sum_{j\neq i}p_{j|i}\log_{2}p_{j|i}.}$
-Intuitively, perplexity controls how large a neighborhood to consider around each point when approximating the topological structure of the data (van der Maaten and Hinton, 2008).
+Intuitively, perplexity controls how large a neighborhood to consider around each point when approximating the topological structure of the data [17].
 As such, it implicitly balances attention to local and global aspects of the data with high values of perplexity placing more emphasis on global aspects.
@@
 $D_{KL}(P||Q)=\sum_{i,j}p_{ij}\log\frac{p_{ij}}{q_{ij}}.$
-The low-dimensional similarities $q_{ij}$ are defined using the t-distribution with one degree of freedom,
+The low-dimensional similarities $q_{ij}$ are defined using the t distribution with one degree of freedom,
 $q_{ij}=\frac{(1+||y_{i}-y_{j}||^{2})^{-1}}{\sum_{k\neq l}(1+||y_{k}-y_{l}||^{2})^{-1}}.$
@@
 2.2 UMAP
-Uniform Manifold Approximation and Projection (UMAP) is another nonlinear dimension reduction method that has been rising in popularity.
+Uniform Manifold Approximation and Projection (UMAP, [12]) is another nonlinear dimension reduction method that has been rising in popularity.
 Like t-SNE, UMAP is a powerful tool for visualizing high-dimensional data that requires user calibration.
-The architecture of the UMAP algorithm is similar to that of t-SNE’s — high-dimensional similarities are computed and the outputted representation is the set of low-dimensional points whose low-dimensional similarities best match the high-dimensional similarities.
+The architecture of the UMAP algorithm is similar to that of t-SNE’s — high-dimensional similarities are computed and the resulting representation is the set of low-dimensional points whose low-dimensional similarities best match the high-dimensional similarities.
 The cost function and the formulas for high/low-dimensional similarities differ from t-SNE.
-See (McInnes et al., 2020) for details.
+See [12] for details.
 UMAP has a hyperparameter called n_neighbors that is analogous to t-SNE’s perplexity.
 It determines how large a neighborhood to use around each point when calculating high-dimensional similarities.
-The original authors make no recommendation for optimal values of n_neighbors, but their implementation defaults to n_neighbors = 15 (McInnes et al., 2020).
+The original authors make no recommendation for optimal values of n_neighbors, but their implementation defaults to n_neighbors = 15 [12].
 3 Methods
@@
 Suppose we have a reconstruction error function $f(D_{1},D_{2})$ that quantifies how well the data set $D_{2}$ represents the data set $D_{1}$.
-Prior works like (Chari & Pachter, 2023), (Espadoto et al., 2021), (Huang et al., 2022), and (Kodak & Berens, 2019) use various reconstruction error functions to quantify performance; only, they study $f(Z+\epsilon,X)$ to measure how well the constructed representation $X$ represents the original data $Z+\epsilon$.
+Prior works like [4], [7], [9], and [10] use various reconstruction error functions to quantify performance; only, they study $f(Z+\epsilon,X)$ to measure how well the constructed representation $X$ represents the original data $Z+\epsilon$.
 We argue it is more appropriate to compare $X$ against the underlying signal $Y$ by examining $f(Y,X)$.
 Prior works in dimension reduction have suggested various quantitative metrics for measuring dimension reduction performance.
-In line with recent discussions of perplexity [(Kodak & Berens, 2019) and (Chari & Pachter, 2023)], we focus on two different metrics — one that measures local performance and one that measures global performance.
-For local performance, we use a nearest-neighbor type metric called trustworthiness (Venna and Kaski, 2006).
-Let $n$ be the sample size and $r(i,j)$ the rank of point $j$ point among the $k$ nearest neighbors of point $i$ in high dimension.
-Let $U_{k}(i)$ denote the set of points among the $k$ nearest neighbors of point $i$ in low dimension, but not high dimension.
+In line with recent discussions of perplexity ([10] and [4]), we focus on two different metrics — one that measures local performance and one that measures global performance.
+For local performance, we use a nearest-neighbor type metric called trustworthiness [18].
+Let $n$ be the sample size and $r(i,j)$ the rank of point $j$ among the $k$ nearest neighbors of point $i$ in high dimension.
+Let $U_{k}(i)$ denote the set of points among the $k$ nearest neighbors of point $i$ in low dimension, but not in high dimension.
 Then
@@
 $f_{trust}(D_{1},D_{2})\approx 1-\frac{2}{mk(2n-3k-1)}\sum_{l=1}^{m}\sum_{j\in U_{k}(i_{l})}\left[r(i_{l},j)-k\right].$
-For global performance, we use Shepard goodness (Espadoto et al., 2021). Shepard goodness is the Kendall correlation (a rank-based correlation) between high and low-dimensional inter-point distances,
+For global performance, we use Shepard goodness [7]. Shepard goodness is the Kendall correlation (a rank-based correlation) between high and low-dimensional inter-point distances,
 $f_{\textrm{Shep}}(D_{1},D_{2})=\sigma_{\textrm{Kendall}}(||z_{i}-z_{j}||^{2},||\varphi(z_{i})-\varphi(z_{j})||^{2}).$
@@
 3.3 Using this framework
-When using our framework to simulate examples, three components must be specified: $Z+\epsilon$, $Y$, and $\textrm{Emb}()$.
+When using our framework to model examples, three components must be specified: $Z+\epsilon$, $Y$, and $\textrm{Emb}()$.
 These elements describe the original data, the underlying signal, and the embedding of the signal in data space, respectively.
@@
 $Z+\epsilon=\begin{bmatrix}Y&|&0\end{bmatrix}+\epsilon.$
-Figure 3a: Shepard Goodness vs. Trustworthiness (Links)
-Figure 3b: Trustworthiness vs. Perplexity (Links)
 Practical examples are more tricky because we do not have the luxury of first defining $Y$.
@@
 where $r$ is the dimension of the projection.
-For a reasonable $r$, we would expect the first $r$ principal components to contain most of the signal, while excluding most of the noise.
+For a reasonable number of dimensions $r$, we would expect the first $r$ principal components to contain most of the signal, while excluding most of the noise.
 Another advantage to using PCA is it gives rise to a natural $\textrm{Emb}()$ function — the PCA inverse transform.
@@
 where $V_{r}\in\mathbb{R}^{p\times r}$ contains the first $r$ eigenvectors of $(Z+\epsilon)^{T}(Z+\epsilon)$ as column vectors.
+Figure 2: Low-Dimensional Simulated Examples; links and trefoil from [20], mammoth from [19]
+Figure 3a: Shepard Goodness vs. Trustworthiness (Links)
 4 Results
 4.1 Simulated Examples
-We first looked at simulated examples with explicitly defined signal structures: inter-linked circles, the trefoil knot, and the mammoth dataset (Figure 2).
-Figure 2: Simulated Examples; links and trefoil from (Wattenberg et al., 2016), mammoth from (Wang et al., 2021)
-Figure 4a: Shepard Goodness vs. Trustworthiness (CyTOF)
-Figure 4b: Trustworthiness vs. Perplexity (CyTOF)
-For each example, the signal $Y$ consisted of 500 points in three dimensions.
+We first looked at simulated examples with explicitly defined signal structures – three low-dimensional examples (Figure 2) and one high-dimensional example.
+For the three low-dimensional examples, the signal $Y$ consisted of 500 points in three dimensions.
 The dataset $Z+\epsilon$ was constructed from $Y$ by adding seven superfluous dimensions and isotropic Gaussian noise to all 10 dimensions.
-We then ran t-SNE using the $R$ package Rtsne with perplexities ranging from 10 to 160333Rtsne implements a Barnes-Hut approximation that only allows for a maximum perplexity of $\frac{n-1}{3}$..
+We then ran t-SNE using the $R$ package Rtsne with perplexities ranging from 10 to 160.
 For each value of perplexity, we ran the algorithm 20 times to mimic the ordinary t-SNE workflow.
@@
 perplexity (Figure 3b).
+Figure 3b: Trustworthiness vs. Perplexity (Links)
+Figure 4: Trustworthiness vs. Perplexity (High-Dimensional Clusters)
 Past a certain perplexity, local performance begins to deteriorate with increasing perplexity, exhibiting the trade-off between global and local performance.
 This cutoff point, however, varies between the two plots.
-When comparing against the original data, a perplexity of 30 maximizes trustworthiness, which is consistent with the original authors’ suggestion of 5 to 50 for perplexity (van der Maaten and Hinton, 2008).
+When comparing against the original data, a perplexity of 30 maximizes trustworthiness, which is consistent with the original authors’ suggestion of 5 to 50 for perplexity [17].
 When comparing against the signal, a perplexity of 110 maximizes trustworthiness.
@@
 Conversely, larger perplexities lead to more stable representations that are less affected by noise.
-The other two simulated examples exhibit a similar trend (See Supplementary Information).
-Figure 5: Trustworthiness vs. Perplexity for $r=5$ (CyTOF)
-4.2 Real Examples
-For a practical example, we looked at a CyTOF dataset containing 239,933 observations in 49 dimensions (Strauss-Albee et al., 2015).
+The other two low-dimensional simulated examples exhibit a similar trend (See Table 1 and Supplementary Information).
+Figure 5: Scree Plot for CyTOF Dataset
+Figure 6a: Shepard Goodness vs. Trustworthiness (CyTOF)
+Figure 6b: Trustworthiness vs. Perplexity (CyTOF)
+The high-dimensional example was constructed by generating seven 10-dimensional Gaussian clusters with varying centers.
+More specifically, the signal $Y$ contained 210 points in 10 dimensions.
+The dataset $Z+\epsilon$ was constructed from $Y$ by adding 50 superfluous dimensions and isotropic Gaussian noise to all 60 dimensions.
+Again, local performance peaked at different perplexities when changing the frame of reference (Figure 4).
+When comparing against the original data, trustworthiness was maximized at a perplexity of 40.
+When comparing against the underlying signal, trustworthiness was maximized at a perplexity of 50.
+4.2 Practical Examples
+In addition to simulated datasets, we looked at three practical datasets: a cytometry by time-of-flight (CyTOF) dataset [15], a single-cell RNA sequencing dataset [16], and a microbiome dataset [2].
+For each dataset, we compared the optimal perplexity for locally replicating the original data versus just the signal at varying numbers of signal dimensions.
+We explore the CyTOF dataset in detail here.
+Summaries for the other two practical examples can be found in Table 1, and the data processing details for the other two practical examples can be found in the Supplementary Information.
+The CyTOF dataset contained 239,933 observations in 49 dimensions [15].
 To reduce the computational load, a subset of 5,000 observations was sampled.
-In line with the ordinary t-SNE workflow, a log transformation was followed by a PCA pre-processing step to reduce the number of dimensions to 10.
-The processed data set to be studied consisted of 5,000 observations in 10 dimensions,
-$Z+\epsilon\in\mathbb{R}^{5,000\times 10}.$
-$Y$ was extracted by taking the first three principal components,
-$Y=\textrm{PCA}_{3}(Z+\epsilon).$
-We computed the t-SNE representations for perplexity values ranging from 20 to 1,600.
+In line with the ordinary t-SNE workflow, a log transformation was followed by a PCA pre-processing step to reduce the number of dimensions to 30, which still retained 77% of the variance in the original data.
+The processed data set to be studied consisted of 5,000 observations in 30 dimensions,
+$Z+\epsilon\in\mathbb{R}^{5,000\times 30}.$
+To determine the dimension of the signal, we drew a scree plot (Figure 5), which depicts flatening after the fifth eigenvalue.
+And thus, $Y$ was extracted by taking the first five principal components,
+$Y=\textrm{PCA}_{5}(Z+\epsilon).$
+We computed the t-SNE representations for perplexity values ranging from 10 to 300.
 For each perplexity value, 20 different t-SNE representations were computed.
@@
 $f_{\textrm{trust}}(Y,X)$.
-As with the simulated examples, there is a difference in trend when the frame of reference is the original data versus when it’s just the underlying signal.
-More specifically, the t-SNE representations better locally represent the original data for lower perplexities and better locally represent the underlying signal for larger perplexities (Figure 4b).
-When compared against the original data, trustworthiness is maximized at a perplexity of 50, which is consistent with (Kobak and Berens, 2019)’s recommendation of setting perplexity to $n/100$.
-When compared against the underlying signal, trustworthiness is maximized at a much larger perplexity of 900, reinforcing the hypothesis that lower values of perplexity may be overfitting the noise.
-If we, instead, decide to be a little more conservative and use the first five principal components to represent the signal, we still see a similar trend (Figure 5).
+As with the simulated examples, there is a difference in trend when the frame of reference is the original data versus when it’s just the underlying signal (Figure 6b).
+When compared against the original data, trustworthiness is maximized at a perplexity of 50, which is consistent with [10]’s recommendation of setting perplexity to $n/100$.
+When compared against the underlying signal, trustworthiness is maximized at a larger perplexity of 220, reinforcing the hypothesis that lower values of perplexity may be overfitting the noise.
+Figure 7: Trustworthiness vs. Perplexity for $r=8$ (CyTOF)
+If we, instead, decide to be a little more conservative and use the first eight principal components to represent the signal, we still see a similar trend (Figure 7).
 Trustworthiness still increases then decreases with perplexity.
-In this case, however, trustworthiness peaks at a perplexity of 100, less than before.
-By including two extra principal components in the signal, we’re assuming the data contains less noise, allowing the model to be more aggressive during the fitting process.
-Further experimentation confirmed that the trustworthiness-maximizing perplexity decreased as we included more dimensions in the signal.
+When compared against the original data, trustworthiness is maximized at a perplexity of 45.
+When compared against the signal, trustworthiness is maximized at a perplexity of 65.
+By including three extra principal components in the signal, we’re assuming the data contains less noise, allowing the model to be more aggressive during the fitting process.
+4.3 Summary of Results
+See Table 1 for a summary of the results.
+$n$, $p$, and $r$ represent the sample size, dimension of the (post PCA-processed) data, and dimension of the extracted signal, respectively.
+The optimal perplexity when comparing against the signal was greater than the optimal perplexity when comparing against the original data for every example.
 4.4 UMAP and n_neighbors
 If n_neighbors functions similarly to perplexity, we’d expect small values of n_neighbors to overfit the data as well.
-Our experiments suggest this to be the case, but to a lesser degree.
-An identical experiment was run using the Python package umap-learn.
-n_neighbor values ranging from 20 to 1,000 were tested on the same CyTOF data set.
-An n_neighbors value of 20 maximized trustworthiness when comparing against the original data, but an n_neighbors value of 600 maximized trustworthiness when comparing against the underlying signal (see Supplementary Information for plots).
-While the difference in local performance for different values of n_neighbors was smaller, the UMAP results exhibited a similar trend.
-Local performance peaked at a larger value when compared against the underlying signal versus the original data.
+An identical experiment was run using the Python package umap-learn on the CyTOF data.
+n_neighbor values ranging from 10 to 300 were tested on the same CyTOF data set.
+An n_neighbors value of 120 maximized trustworthiness when comparing against the original data, but an n_neighbors value of 160 maximized trustworthiness when comparing against the underlying signal.
+See Supplementary Information for plots.
 5 Application
@@
 The signal should include the features of the data one desires to retain throughout the dimension reduction process.
-When using a PCA projection to serve as the signal, one could draw a scree plot or employ a component selection algorithm such as parallel analysis to determine the dimension of the signal (Horn, 1965).
+When using a PCA projection to serve as the signal, one could draw a scree plot or employ a component selection algorithm such as parallel analysis [8] to determine the dimension of the signal.
 With a signal constructed, it remains to compute t-SNE outputs at varying perplexities.
@@
 From there, one can choose the representation with the desirable balance of local and global performance.
-See algorithm 1.
+A summary is given in Algorithm 1.
 Sample code is available at https://github.com/JustinMLin/DR-Framework/.
+Optimal Perplexity
+Dataset $n$ $p$ $r$ signal + noise signal
+Links [20] 500 10 3 30 110
+Trefoil [20] 500 10 3 30 110
+Mammoth [19] 500 10 3 20 100
+High-Dimensional Clusters 210 60 10 40 50
+CyTOF [15] 5,000 30 5 50 220
+CyTOF [15] 5,000 30 8 45 65
+scRNA-seq [16] 864 500 5 40 120
+scRNA-seq [16] 864 500 10 50 60
+Microbiome [2] 280 66 5 50 90
+Microbiome [2] 280 66 8 60 85
+Table 1: Summary of results
+It is worth noting that computational barriers may arise, especially for very large data sets.
+To alleviate such issues, trustworthiness and Shepard goodness can be approximated by subsampling before calculation.
+Furthermore, t-SNE is generally robust to small changes in perplexity [17], so checking a handful of different perplexities is sufficient.
+If computing the t-SNE representations is the limiting factor, the perplexity can be calibrated for a subsample of the data, instead.
+[14] found that embedding a $\rho$-sample, where $\rho\in(0,1]$ is the sampling rate, with perplexity $\textrm{Perp}^{\prime}$ gives a visual impression of embedding the original data with perplexity
+$\textrm{{Perp}}=\frac{\textrm{Perp}^{\prime}}{\rho}.$
+With these concessions, applying this framework to calibrate perplexity is feasible for data sets of any size.
 Algorithm 1 Measuring Performance in the Presence of Noise
@@
 12:Choose output with desired balance of local and global performance
-It is worth noting that computational barriers may arise, especially for very large data sets.
-To alleviate such issues, trustworthiness and Shepard goodness can be approximated by subsampling before calculation.
-Furthermore, t-SNE is generally robust to small changes in perplexity (van der Maaten and Hinton, 2008), so checking a handful of different perplexities is sufficient.
-If computing the t-SNE representations is the limiting factor, the perplexity can be calibrated for a subsample of the data, instead.
-(Skrodzki et al., 2023) found that embedding a $\rho$-sample, where $\rho\in(0,1]$ is the sampling rate, with perplexity $\textrm{Perp}^{\prime}$ gives a visual impression of embedding the original data with perplexity
-$\textrm{{Perp}}=\frac{\textrm{Perp}^{\prime}}{\rho}.$
-With these concessions, applying this framework to calibrate perplexity is feasible for data sets of any size.
 6 Discussion
@@
 Low perplexities lead to overly-flexible t-SNE models that are heavily impacted by the presence of noise, while higher perplexities exhibit better performance due to increased stability.
-These considerations are especially important when working with heavily noised data, which are especially prevalent in the world of single-cell transcriptomics (Chu et al., 2022).
+These considerations are especially important when working with heavily noised data, which are especially prevalent in the world of single-cell transcriptomics [5].
 We have also presented a framework for modeling dimension reduction problems in the presence of noise.
@@
 7 Data Availability
-All data and example code are freely available at https://github.com/JustinMLin/DR-Framework/.
+All data and code are freely available at https://github.com/JustinMLin/DR-Framework/.
+Supplementary Information
+SI Simulated Examples
+SI.1 Trefoil Plots
+Figure S1: Trefoil Plots
+SI.2 Mammoth Plots
+Figure S2: Mammoth Plots
+SII Practical Examples
+SII.1 UMAP Plots (CyTOF)
+Figure S3: UMAP Plots
+SII.2 scRNA Dataset
+This is a dataset of induced pluripotent stem cells generated from three different individuals [16].
+The original data includes 864 units and 19,027 readings per unit.
+To process this zero-inflated count data, columns containing a large proportion of 0’s (20% or more) were removed before a log transformation was applied.
+This reduced the dimension to 5,431.
+A PCA pre-processing stop further reduced the dimension to 500, which still retained 88% of the variance of the log-transformed data.
+The signal was first taken to be the first five principal components, then the first 10 principal components.
+Notice the optimal perplexity when compared against the original data differed between these two experiments, even though it should theoretically be independent of the chosen signal dimension.
+This is due to the inherent randomness of the t-SNE algorithm.
+Figure S4: scRNA Plots (r = 5)
+Figure S5: scRNA Plots (r = 10)
+SII.3 Microbiome Dataset
+[2] compares the faecal microbial communities from 22 subjects using complete shotgun DNA sequencing.
+The original data contained 280 samples and 553 genera.
+To deal with a large number of near-zero readings, columns containing a large proportion of values less than $10^{-6}$ (60% or more) were removed.
+This reduced the dimension to 66.
+A PCA pre-processing was used to center and re-scale the data.
+The signal was first taken to be the first five principal components, then the first eight principal components.
+Notice the optimal perplexity when compared against the original data differed between these two experiments, even though it should theoretically be independent of the chosen signal dimension.
+This is due to the inherent randomness of the t-SNE algorithm.
+Figure S6: Microbiome Plots (r = 5)
+Figure S7: Microbiome Plots (r = 8)

### v5

@@
 These tools are employed for a variety of reasons such as noise reduction, visualization, and to lower computational costs.
-However, there is a fundamental issue that is highly discussed in other modeling problems, but almost entirely ignored in the dimension reduction literature: overfitting.
-If we interpret data as a combination of signal and noise, prior works judge dimension reduction techniques on their ability to capture the entirety of the data, i.e.
-both the signal and the noise.
-In the context of other modeling problems, techniques such as feature-selection, cross-validation, and regularization are employed to combat overfitting, but no such precautions are taken when performing dimension reduction.
-In this paper, we present a framework that models dimension reduction problems in the presence of noise and use this framework to explore the role perplexity and number of neighbors play in overfitting data when applying t-SNE and UMAP.
-More specifically, we show previously recommended values for perplexity and number of neighbors are too small and tend to overfit the noise.
-We also present a workflow others may use to calibrate hyperparameters in the presence of noise.
+However, there is a fundamental issue that is discussed in other modeling problems that is often overlooked in dimension reduction — overfitting.
+In the context of other modeling problems, techniques such as feature-selection, cross-validation, and regularization are employed to combat overfitting, but rarely are such precautions taken when applying dimension reduction.
+Prior applications of the two most popular non-linear dimension reduction methods, t-SNE and UMAP, fail to acknowledge data as a combination of signal and noise when assessing performance.
+These methods are typically calibrated to capture the entirety of the data, not just the signal.
+In this paper, we demonstrate the importance of acknowledging noise when calibrating hyperparameters and present a framework that enables users to do so.
+We use this framework to explore the role hyperparameter calibration plays in overfitting the data when applying t-SNE and UMAP.
+More specifically, we show previously recommended values for perplexity and n_neighbors are too small and overfit the noise.
+We also provide a workflow others may use to calibrate hyperparameters in the presence of noise.
 1 Introduction
 In recent years, non-linear dimension reduction techniques have been growing in popularity due to their usefulness when analyzing high-dimensional data.
-To date, they are the most effective tools for visualizing high-dimensional data.
-Visualization plays a key role in identifying and understanding relationships within the data.
-It exposes patterns, trends, and outliers that can’t be discerned from the raw data.
-The most popular non-linear dimension reduction technique is t-distributed stochastic neighbor embedding (t-SNE, [17]).
-Since the introduction of t-SNE, hyperparameter calibration has proven to be a difficult task.
-The most crucial hyperparameter, known as perplexity, controls how large a neighborhood to consider around a point when determining its location.
-Perplexity calibration is so troublesome, that perplexity-free versions of t-SNE have been proposed [6].
-It is also an extremely important task, since t-SNE is known to produce unfaithful results when mishandled [9].
-The original authors suggested values between 5 and 50 [17], while recent works have suggested perplexities as large as one percent of the sample size [10].
-[3] studied the inverse relationship between perplexity and Kullback-Leibler divergence to design an automatic calibration process that “generally agrees with experts’ consensus.” Manual tuning of perplexity requires a deep understanding of the t-SNE algorithm and general knowledge of the data’s structure.
-When the data’s structure is available, we can visualize the results and choose the perplexity that best captures the hypothesized structure.
+Biologists use these techniques for a variety of visualization and analytic purposes, including exposing cell subtypes [1], checking for batch effects [2], and visualizing the trajectories of differentiating cells [3].
+The most popular non-linear dimension reduction methods are t-distributed Stochastic Neighbor Embedding (t-SNE, [4]) and Uniform Manifold Approximation and Projection (UMAP, [5]).
+Both methods have been applied to various types of data within biology ([1], [6], [7]).
+Since the introduction of t-SNE and UMAP, hyperparameter calibration has proven to be a difficult task.
+The most crucial hyperparameters, t-SNE’s perplexity and UMAP’s n_neighbors, control how large a neighborhood to consider around each point when determining its location in low dimension.
+Calibration is so troublesome, that perplexity-free versions of t-SNE have been proposed [8].
+It is also an extremely important task, since both methods are known to produce unfaithful results when mishandled [9].
+For t-SNE, the original authors suggested perplexities between 5 and 50 [4], while recent works have suggested perplexities as large as one percent of the sample size [10].
+[11] studied the inverse relationship between perplexity and Kullback-Leibler divergence to design an automatic calibration process that “generally agrees with experts’ consensus.” For UMAP, the original authors make no recommendation for optimal values of n_neighbors, but their implementation defaults to n_neighbors = 15 [5].
+Manual tuning of perplexity and n_neighbors requires a deep understanding of the t-SNE and UMAP algorithms, as well as a general knowledge of the data’s structure.
+The primary purpose of dimension reduction is to simplify data in a way that eliminates superfluous or nonessential information, i.e.
+noise.
+Each dimension reduction method does this slightly differently, but most require hyperparameter calibration.
+For example, the classical linear method, PCA, requires tuning of the number of principal components.
+A more contemporary method in biology, PHATE (Potential of Heat-diffusion for Affinity-based Trajectory Embedding) [3], requires tuning of a hyperparameter named diffusion time scale $t$.
+PHATE represents the structure of the data by computing local similarities then walking through the data using a Markovian random-walk diffusion process.
+$t$ determines the number of steps taken in a random walk and “provides a tradeoff between encoding local and global information in the embedding” [3].
+Perplexity and n_neighbors serve the same purpose in their respective algorithms.
+Hence, we believe t-SNE and UMAP are capable of handling noise, but naïve calibrations that disregard noise often result in overfitting.
+To assess dimension reduction performance in the presence of noise, we must acknowledge noise during the evaluation process.
+When the data’s structure is available, we can visualize the results and choose the representation that best captures the hypothesized structure.
 In supervised problems, for example, we look for low-dimensional representations that cluster according to the class labels.
 For unsupervised problems, however, the structure is often unknown, so we cannot visually assess each representation.
-In these cases, we must resort to quantitative measures of performance to understand how the well the low-dimensional representation represents the high-dimensional data.
-While this strategy is heavily discussed in the machine learning literature, prior works disregard the possibility of overfitting when measuring performance.
+In these cases, we must resort to quantitative measures of performance to understand how well the low-dimensional representation reproduces the high-dimensional data.
+While this strategy is heavily discussed in the machine learning literature, many prior works disregard the possibility of overfitting when quantitatively measuring performance.
 In this paper, we present a framework for studying dimension reduction methods in the presence of noise (Section 3).
 We then use this framework to calibrate t-SNE and UMAP hyperparameters in both simulated and practical examples to illustrate how the disregard of noise leads to miscalibration (Section 4).
-We also discuss how other researchers may use this framework in their own work (Section 5).
+We also discuss how other researchers may use this framework in their own work (Section 5) and present a case study that walks the reader through the application of the framework to a modern data set (Section 6).
 2 Background
@@
 $p_{j|i}=\frac{\exp(-||x_{i}-x_{j}||^{2}/2\sigma_{i}^{2})}{\sum_{k\neq i}\exp(-||x_{i}-x_{k}||^{2}/2\sigma_{i}^{2})}.$
-Thus for each point $x_{i}$, we have a probability distribution $P_{i}$ that quantifies the similarity of every other point to $x_{i}$.
+Thus for each point $x_{i}$, we have a probability distribution $P_{i}$ that quantifies the similarity of $x_{i}$ to every other point.
 The scale of the Gaussian kernel, $\sigma_{i}$, is chosen so that the perplexity of the probability distribution $P_{i}$, in the information theory sense, is equal to a pre-specified value also named perplexity,
 $\textrm{perplexity}=2^{-\sum_{j\neq i}p_{j|i}\log_{2}p_{j|i}.}$
-Intuitively, perplexity controls how large a neighborhood to consider around each point when approximating the topological structure of the data [17].
+Intuitively, perplexity controls how large a neighborhood to consider around each point when approximating the topological structure of the data.
 As such, it implicitly balances attention to local and global aspects of the data with high values of perplexity placing more emphasis on global aspects.
@@
 $q_{ij}=\frac{(1+||y_{i}-y_{j}||^{2})^{-1}}{\sum_{k\neq l}(1+||y_{k}-y_{l}||^{2})^{-1}}.$
-The main downsides of t-SNE are its inherit randomness and sensitivity to hyperparameter calibration.
-The minimization of the KL divergence is done using gradient descent methods with incorporated randomness to avoid stagnating at local minima.
+The primary downsides of t-SNE are its inherent randomness, unintuitive results, and sensitivity to hyperparameter calibration.
+The minimization of KL divergence is done using gradient descent methods with incorporated randomness to avoid stagnating at local minima.
 As a result, the output differs between runs of the algorithm.
 Hence, the traditional t-SNE workflow often includes running the algorithm multiple times at various perplexities before choosing the best representation.
+t-SNE is also known to produce results that are not faithful to the true structure of the data, even when calibrated correctly.
+For example, cluster sizes and inter-cluster distances aren’t always consistent with the original data [13].
+Such artifacts of the t-SNE algorithm can be confused for significant structures by inexperienced users.
 2.2 UMAP
 Uniform Manifold Approximation and Projection (UMAP, [5]) is another nonlinear dimension reduction method that has been rising in popularity.
-Like t-SNE, UMAP is a powerful tool for visualizing high-dimensional data that requires user calibration.
-The architecture of the UMAP algorithm is similar to that of t-SNE’s — high-dimensional similarities are computed and the resulting representation is the set of low-dimensional points whose low-dimensional similarities best match the high-dimensional similarities.
-The cost function and the formulas for high/low-dimensional similarities differ from t-SNE.
+Originally introduced as a more computationally efficient alternative to t-SNE, UMAP is a powerful tool for visualizing high-dimensional data that requires user calibration.
+While its underlying ideology is completely different from that of t-SNE, the UMAP algorithm is very similar architecturally to the t-SNE algorithm — high-dimensional similarities are computed and the resulting representation is the set of low-dimensional points whose low-dimensional similarities best match the high-dimensional similarities.
 See [5] for details.
-UMAP has a hyperparameter called n_neighbors that is analogous to t-SNE’s perplexity.
-It determines how large a neighborhood to use around each point when calculating high-dimensional similarities.
-The original authors make no recommendation for optimal values of n_neighbors, but their implementation defaults to n_neighbors = 15 [12].
+The largest difference is UMAP’s default initialization process.
+UMAP uses Laplacian eigenmaps to initialize the low-dimensional representation, which is then adjusted to minimize the cost function.
+Most t-SNE implementations use PCA during the initialization process.
+The initialization process is the primary benefit of the default implementation of UMAP, but t-SNE and UMAP have been shown to perform similarly with identical initializations [7].
+Modern implementations of both algorithms are also comparable in speed.
+UMAP shares similar disadvantages with t-SNE. It can create unfaithful representations that require experience to interpret and is sensitive to hyperparameter calibration [14].
 3 Methods
@@
 Prior works quantitatively measure how well low-dimensional representations match the high-dimensional data.
-However, if we view the original data as a composition of signal and noise, we must not reward capturing the noise.
-This would lead to overfitting.
+However, if we consider data as a composition of signal and noise, we must not reward capturing the noise.
 Therefore, we should be comparing the low-dimensional representation against the signal underlying our data, rather than the entirety of the data.
-Suppose the underlying signal of our data is captured by an $r$-dimensional matrix $Y\in\mathbb{R}^{n\times r}$.
-In the context of dimension reduction, the underlying signal is often lower dimension than the original data.
+Suppose the underlying signal of our data is described by an $r$-dimensional matrix $Y\in\mathbb{R}^{n\times r}$.
+In the context of dimension reduction, the signal is often lower dimension than the original data.
 Let $p\geq r$ be the dimension of the original data set, and let $\textrm{Emb}:\mathbb{R}^{r}\to\mathbb{R}^{p}$ be the function that embeds the signal in data space.
@@
 We then assume the presence of random error.
-The original data can then be modeled by
-$Z+\epsilon\textrm{ for }\epsilon\sim N_{p}(0,\Sigma).$
-The dimension reduction method $\varphi$ is applied to $Z+\epsilon$ to get a low-dimensional representation $X\in\mathbb{R}^{n\times q}$. See Figure 1.
-{tikzcd}
+The original data can then be modeled by $Z+\epsilon\textrm{ for }\epsilon\sim N_{p}(0,\Sigma)$.
+The dimension reduction method $\varphi$ is applied to $Z+\epsilon$ to get a low-dimensional representation $X\in\mathbb{R}^{n\times q}$.
+See Figure 1.
+${Z+\epsilon}$${X}$${Y}$${Z}$$\scriptstyle{\varphi}$EmbDimension (low to high)
 Figure 1: Dimension reduction framework
@@
 Prior works like [9], [10] , [15], and [16] use various reconstruction error functions to quantify performance; only, they study $f(Z+\epsilon,X)$ to measure how well the constructed representation $X$ represents the original data $Z+\epsilon$.
-We argue it is more appropriate to compare $X$ against the underlying signal $Y$ by examining $f(Y,X)$.
+We argue it is more appropriate to compare $X$ against the signal $Y$ by examining $f(Y,X)$.
 Prior works in dimension reduction have suggested various quantitative metrics for measuring dimension reduction performance.
-In line with recent discussions of perplexity ([10] and [4]), we focus on two different metrics — one that measures local performance and one that measures global performance.
+In line with recent discussions of perplexity ([10] and [15]), we employ two different metrics — one that measures local performance and one that measures global performance.
 For local performance, we use a nearest-neighbor type metric called trustworthiness [17].
@@
 $f_{trust}(D_{1},D_{2})\approx 1-\frac{2}{mk(2n-3k-1)}\sum_{l=1}^{m}\sum_{j\in U_{k}(i_{l})}\left[r(i_{l},j)-k\right].$
-For global performance, we use Shepard goodness [7]. Shepard goodness is the Kendall correlation (a rank-based correlation) between high and low-dimensional inter-point distances,
-$f_{\textrm{Shep}}(D_{1},D_{2})=\sigma_{\textrm{Kendall}}(||z_{i}-z_{j}||^{2},||\varphi(z_{i})-\varphi(z_{j})||^{2}).$
+Local performance is the primary concern when applying t-SNE and UMAP, so our experiments focus on maximizing trustworthiness.
+For global performance, we use Shepard goodness [16]. Shepard goodness is the Spearman correlation, a rank-based correlation, between high and low-dimensional inter-point distances,
+$f_{\textrm{Shep}}(D_{1},D_{2})=\sigma_{\textrm{Spearman}}(||z_{i}-z_{j}||,||\varphi(z_{i})-\varphi(z_{j})||).$
 Again for very large values of $n$, we can get an approximation by calculating the correlation between inter-point distances of a random subsample.
 3.3 Using this framework
-When using our framework to model examples, three components must be specified: $Z+\epsilon$, $Y$, and $\textrm{Emb}()$.
+When using this framework to model examples, three components must be specified: $Z+\epsilon$, $Y$, and $\textrm{Emb}()$.
 These elements describe the original data, the underlying signal, and the embedding of the signal in data space, respectively.
 When simulating examples, it’s natural to start with the underlying signal $Y$ then construct $Z+\epsilon$ by attaching extra dimensions and adding Gaussian noise.
-The $\textrm{Emb}()$ function is then given by
-$\textrm{Emb}(y)=(y,0,\ldots,0),$
-so that
+The $\textrm{Emb}()$ function is then given by $\textrm{Emb}(y)=(y,0,\ldots,0)$ so that
 $Z+\epsilon=\begin{bmatrix}Y&|&0\end{bmatrix}+\epsilon.$
+Figure 2: Low-Dimensional Simulated Examples; links and trefoil from [13], mammoth from [18]
 Practical examples are more tricky because we do not have the luxury of first defining $Y$.
 Instead, we are given the data $Z+\epsilon$ from which we must extract $Y$, or at least our best estimate.
-This process is left to the researcher and should be based on a priori knowledge of the data.
+This process is dependent on the specifics of the problem and should be based on a priori knowledge of the data.
 If there is no specific signal of interest, a more general approach can be taken.
-We used a PCA projection of the data to represent the signal,
-$Y=\textrm{PCA}_{r}(Z+\epsilon),$
-where $r$ is the dimension of the projection.
-For a reasonable number of dimensions $r$, we would expect the first $r$ principal components to contain most of the signal, while excluding most of the noise.
+We used a PCA projection of the data to represent the signal, $Y=\textrm{PCA}_{r}(Z+\epsilon)$, where $r$ is the dimension of the projection.
+For a reasonably chosen $r$, we expect the first $r$ principal components to contain most of the signal, while excluding most of the noise.
 Another advantage to using PCA is it gives rise to a natural $\textrm{Emb}()$ function — the PCA inverse transform.
@@
 where $V_{r}\in\mathbb{R}^{p\times r}$ contains the first $r$ eigenvectors of $(Z+\epsilon)^{T}(Z+\epsilon)$ as column vectors.
-Figure 2: Low-Dimensional Simulated Examples; links and trefoil from [20], mammoth from [19]
-Figure 3a: Shepard Goodness vs. Trustworthiness (Links)
+Figure 3: Trustworthiness vs. Perplexity (Links $sd=1$)
+Figure 4: Trustworthiness-Maximizing Representations (Links $sd=1$)
 4 Results
@@
 We first looked at simulated examples with explicitly defined signal structures – three low-dimensional examples (Figure 2) and one high-dimensional example.
-For the three low-dimensional examples, the signal $Y$ consisted of 500 points in three dimensions.
-The dataset $Z+\epsilon$ was constructed from $Y$ by adding seven superfluous dimensions and isotropic Gaussian noise to all 10 dimensions.
-We then ran t-SNE using the $R$ package Rtsne with perplexities ranging from 10 to 160.
-For each value of perplexity, we ran the algorithm 20 times to mimic the ordinary t-SNE workflow.
-If we were to disregard the distinction between signal and noise, a plot of $f_{\textrm{Shep}}(Z+\epsilon,X)$ vs.
-$f_{\textrm{trust}}(Z+\epsilon,X)$ could be used to calibrate perplexity.
-To avoid overfitting the noise, a plot of $f_{\textrm{Shep}}(Y,X)$ vs.
-$f_{\textrm{trust}}(Y,X)$ should be used instead.
-See Figure 3a for plots of the links example.
-Both plots depict an increase in global performance as perplexity increases.
-Both plots also depict an increase in local performance followed by a decrease as perplexity increases.
-This trend is more evident when we plot trustworthiness vs.
-perplexity (Figure 3b).
-Figure 3b: Trustworthiness vs. Perplexity (Links)
-Figure 4: Trustworthiness vs. Perplexity (High-Dimensional Clusters)
-Past a certain perplexity, local performance begins to deteriorate with increasing perplexity, exhibiting the trade-off between global and local performance.
+The links example and the high-dimensional example are explored here.
+See Table 1 and the Supporting Information (SI.1 and SI.2) for the other simulated examples.
+4.1.1 Links Data Set
+For the links example, the signal $Y$ consisted of two interlocked circles, each containing 250 points, embedded in three dimensions.
+$Z+\epsilon$ was constructed by adding seven superfluous dimensions and isotropic Gaussian noise.
+Various degrees of noise were tested ($sd=0.5,1,1.5,2,2.5,3$).
+t-SNE was run using the $R$ package Rtsne [19] at varying perplexities.
+For each perplexity, the algorithm was run 40 times to mimic the ordinary t-SNE workflow.
+If the distinction between signal and noise were disregarded, a plot of $f_{\textrm{trust}}(Z+\epsilon,X)$ vs.
+perplexity could be used to maximize local performance.
+To avoid overfitting the noise, a plot of $f_{\textrm{trust}}(Y,X)$ vs.
+perplexity should be used instead.
+See Figure 3 for examples of these plots for the $sd=1$ case.
+Both plots depict an increase in local performance followed by a decrease as perplexity increases.
 This cutoff point, however, varies between the two plots.
-When comparing against the original data, a perplexity of 30 maximizes trustworthiness, which is consistent with the original authors’ suggestion of 5 to 50 for perplexity [17].
-When comparing against the signal, a perplexity of 110 maximizes trustworthiness.
-We hypothesize t-SNE tends to overfit the noise when the perplexity is too low.
+When comparing against the original data, the trustworthiness-maximizing representation was constructed with a perplexity of 40, which is consistent with the original authors’ suggestion of 5 to 50 for perplexity [4].
+When comparing against the signal, the trustworthiness-maximizing representation was constructed with a perplexity of 80.
+With the signal structure known, we are also able to visually assess the trustworthiness-maximizing representations.
+Figure 4 shows the trustworthiness-maximizing representations for the $sd=1$ case.
+Notice the larger perplexity was able to successfully separate the circles in the presence of noise, while the smaller perplexity was not.
+By using the signal as the frame of reference, our framework correctly rewarded the representation that was able to successfully separate the two links.
+The same pattern held true for other levels of noise. The optimal perplexity was consistently larger when comparing against the signal, rather than the original data (Figure 5).
+Figure 5: Optimal Perplexity (Links)
+These results suggest larger perplexities perform better in the presence of noise, both quantitatively and qualitatively.
+We hypothesize t-SNE tends to overfit the noise when the perplexity is too small.
 Intuitively, small perplexities are more affected by slight perturbations of the data when only considering small neighborhoods around each point, leading to unstable representations.
-Conversely, larger perplexities lead to more stable representations that are less affected by noise.
-The other two low-dimensional simulated examples exhibit a similar trend (See Table 1 and Supplementary Information).
-Figure 5: Scree Plot for CyTOF Dataset
-Figure 6a: Shepard Goodness vs. Trustworthiness (CyTOF)
-Figure 6b: Trustworthiness vs. Perplexity (CyTOF)
-The high-dimensional example was constructed by generating seven 10-dimensional Gaussian clusters with varying centers.
-More specifically, the signal $Y$ contained 210 points in 10 dimensions.
-The dataset $Z+\epsilon$ was constructed from $Y$ by adding 50 superfluous dimensions and isotropic Gaussian noise to all 60 dimensions.
-Again, local performance peaked at different perplexities when changing the frame of reference (Figure 4).
-When comparing against the original data, trustworthiness was maximized at a perplexity of 40.
-When comparing against the underlying signal, trustworthiness was maximized at a perplexity of 50.
+Conversely, larger perplexities lead to more stable representations that are more robust to noise.
+Figure 6: Trustworthiness vs. Perplexity (High-Dimensional Clusters $sd=3$)
+Figure 7: Trustworthiness-Maximizing Representations (High-Dimensional Clusters $sd=3$)
+Figure 8: Optimal Perplexity (High-Dimensional Clusters)
+Figure 9: Scree Plot for scRNA-seq Data Set
+Figure 10: Trustworthiness vs. Perplexity for r = 5 (scRNA-seq)
+Figure 11: Trustworthiness vs. Perplexity for r = 5 (scRNA-seq)
+4.1.2 High-Dimensional Clusters
+The signal $Y$ consisted of seven Gaussian clusters, each containing 50 points, in seven dimensions.
+The clusters were drawn from multivariate normal distributions with mean $10e_{i}$ and random diagonal covariance matrices, where $e_{i}$ is the $i^{\textrm{th}}$ standard basis vector.
+The data set $Z+\epsilon$ was constructed from $Y$ by adding 53 superfluous dimensions and isotropic Gaussian noise to all 60 dimensions.
+Various degrees of noise were tested $(sd=2,2.5,3,3.5,4,4.5)$.
+When $sd=3$, local performance peaked at different perplexities when changing the frame of reference (Figure 6).
+When comparing against the original data, trustworthiness was maximized at a perplexity of 55.
+When comparing against the signal, trustworthiness was maximized at a perplexity of 60.
+See Figure 7 for the trustworthiness-maximizing representations.
+Visually, both representations maintain the original clustering to some extent, but the higher-perplexity representation shows less mixing between the clusters and had a larger average silhouette width (0.178) than the lower-perplexity representation (0.121).
+This suggests the higher-perplexity representation better maintained the original clustering.
+Figure 8 shows the optimal perplexities for different levels of noise. Again, the trustworthiness-maximizing perplexity was larger when comparing against the signal for all levels of noise.
 4.2 Practical Examples
-In addition to simulated datasets, we looked at three practical datasets: a cytometry by time-of-flight (CyTOF) dataset [15], a single-cell RNA sequencing dataset [16], and a microbiome dataset [2].
-For each dataset, we compared the optimal perplexity for locally replicating the original data versus just the signal at varying numbers of signal dimensions.
-We explore the CyTOF dataset in detail here.
-Summaries for the other two practical examples can be found in Table 1, and the data processing details for the other two practical examples can be found in the Supplementary Information.
-The CyTOF dataset contained 239,933 observations in 49 dimensions [15].
-To reduce the computational load, a subset of 5,000 observations was sampled.
-In line with the ordinary t-SNE workflow, a log transformation was followed by a PCA pre-processing step to reduce the number of dimensions to 30, which still retained 77% of the variance in the original data.
-The processed data set to be studied consisted of 5,000 observations in 30 dimensions,
-$Z+\epsilon\in\mathbb{R}^{5,000\times 30}.$
-To determine the dimension of the signal, we drew a scree plot (Figure 5), which depicts flatening after the fifth eigenvalue.
-And thus, $Y$ was extracted by taking the first five principal components,
-$Y=\textrm{PCA}_{5}(Z+\epsilon).$
-We computed the t-SNE representations for perplexity values ranging from 10 to 300.
-For each perplexity value, 20 different t-SNE representations were computed.
-Figure 6a contains the plots for $f_{\textrm{Shep}}(Z+\epsilon,X)$ vs.
-$f_{\textrm{trust}}(Z+\epsilon,X)$ and $f_{\textrm{Shep}}(Y,X)$ vs.
-$f_{\textrm{trust}}(Y,X)$.
-As with the simulated examples, there is a difference in trend when the frame of reference is the original data versus when it’s just the underlying signal (Figure 6b).
-When compared against the original data, trustworthiness is maximized at a perplexity of 50, which is consistent with [10]’s recommendation of setting perplexity to $n/100$.
-When compared against the underlying signal, trustworthiness is maximized at a larger perplexity of 220, reinforcing the hypothesis that lower values of perplexity may be overfitting the noise.
-Figure 7: Trustworthiness vs. Perplexity for $r=8$ (CyTOF)
-If we, instead, decide to be a little more conservative and use the first eight principal components to represent the signal, we still see a similar trend (Figure 7).
+In addition to simulated data sets, we looked at three practical data sets: a single-cell RNA sequencing data set [20], a cytometry by time-of-flight (CyTOF) data set [21], and a microbiome data set [22].
+For each data set, we compared the optimal perplexity for locally replicating the original data versus the estimated signal.
+We explore the scRNA-seq data set in detail here.
+The results of the other two practical examples can be found in Table 1.
+The details can be found in the Supporting Information (SI.1 and SI.2).
+The scRNA-seq data set was generated from induced pluripotent stem cells collected from three different individuals.
+The original data includes 864 units and 19,027 readings per unit.
+To process this zero-inflated count data, columns containing a large proportion of 0’s (20% or more) were removed before a log transformation was applied.
+This step reduced the number of dimensions to 5,431.
+A PCA pre-processing step further reduced the number of dimensions to 500, which still retained 88% of the variance of the log-transformed data.
+Hence, the processed data set consisted of 864 observations in 500 dimensions, $Z+\epsilon\in\mathbb{R}^{864\times 500}$.
+To determine the dimensionality of the signal, we drew a scree plot (Figure 9).
+Note, the first eigenvalue (2359.357) was cut to fit the plot.
+A conservative estimate is five dimensions, so $Y$ was extracted by taking the first five principal components, $Y=\textrm{PCA}_{5}(Z+\epsilon)$.
+We computed the t-SNE representations for perplexities ranging from 10 to 280.
+For each perplexity, 20 different t-SNE representations were computed.
+As with the simulated examples, there is a difference in trend when switching the frame of reference (Figure 10).
+When compared against the original data, trustworthiness is maximized at a perplexity of 40, which is consistent with [4]’s recommendation of 5 to 50.
+When compared against the signal, trustworthiness is maximized at a larger perplexity of 120, reinforcing the hypothesis that lower values of perplexity may be overfitting the noise.
+Visual inspection of the trustworthiness-maximizing representations reveals the effect of increasing the perplexity (Figure 11).
+A hierarchical clustering of the high-dimensional data was computed, then projected onto the trustworthiness-maximizing representations.
+Both representations depict a similar a structure, but the relative positioning of clusters differs.
+For example, the Class 1 cluster is the rightmost cluster in the perplexity = 40 representation, while the Class 5 cluster is the rightmost cluster in the perplexity = 120 representation.
+Furthermore, the left-to-right order of the Class 3, Class 8, and Class 9 clusters is reversed in both representations.
+In terms of local structure, the Class 3 and Class 7 clusters are better preserved in the perplexity = 40 representation, while the Class 12 cluster is better preserved in the perplexity = 120 representation.
+The perplexity = 40 representation also suggests the Class 2 cluster could potentially contain two separate clusters, but this is not consistent with the high-dimensional data according to the dendrogram and higher-order clusterings.
+The perplexity = 120 representation does not mislead in this way.
+Overall, the lower perplexity leads to tighter-knit clusters as expected.
+However, further investigation reveals the over-clustering may be unfaithful to the original data.
+      Optimal Perplexity
+      Data Set       $n$       $p$       $r$       signal + noise       signal
+      Links [13]       500       10       3       40       80
+      Trefoil [13]       500       10       3       35       100
+      Mammoth [18]       500       10       3       30       80
+      High-Dimensional Clusters       210       60       10       55       60
+      scRNA-seq [20]       864       500       5       40       120
+      scRNA-seq [20]       864       500       10       50       60
+      CyTOF [21]       5,000       30       5       50       110
+      CyTOF [21]       5,000       30       8       45       65
+      Microbiome [22]       280       66       5       50       90
+      Microbiome [22]       280       66       8       60       85
+Table 1: Summary of Results
+Figure 12: Trustworthiness vs. Perplexity for r = 10 (scRNA-seq)
+Figure 13: Trustworthiness vs. n_neighbors for UMAP (scRNA-seq)
+If we, instead, decide to be more conservative and use the first 10 principal components to represent the signal, we still see a similar trend (Figure 12).
 Trustworthiness still increases then decreases with perplexity.
-When compared against the original data, trustworthiness is maximized at a perplexity of 45.
-When compared against the signal, trustworthiness is maximized at a perplexity of 65.
+When compared against the original data, trustworthiness is maximized at a perplexity of 50 (Note the optimal perplexity when compared against the original data differed between the two experiments, even though it should theoretically be independent of the chosen signal dimension.
+This is due to the inherent randomness of the t-SNE algorithm).
+When compared against the signal, trustworthiness is maximized at a perplexity of 60.
 By including three extra principal components in the signal, we’re assuming the data contains less noise, allowing the model to be more aggressive during the fitting process.
@@
 If n_neighbors functions similarly to perplexity, we’d expect small values of n_neighbors to overfit the data as well.
-An identical experiment was run using the Python package umap-learn on the CyTOF data.
-n_neighbor values ranging from 10 to 300 were tested on the same CyTOF data set.
-An n_neighbors value of 120 maximized trustworthiness when comparing against the original data, but an n_neighbors value of 160 maximized trustworthiness when comparing against the underlying signal.
-See Supplementary Information for plots.
+An identical experiment was run using the Python package umap-learn [5] on the scRNA-seq data.
+n_neighbor values ranging from 10 to 300 were tested and an n_neighbors value of 190 maximized trustworthiness when comparing against the original data, but an n_neighbors value of 300 maximized trustworthiness when comparing against the signal (Figure 13).
 5 Application
@@
 When using a PCA projection to serve as the signal, one could draw a scree plot or employ a component selection algorithm such as parallel analysis [23] to determine the dimension of the signal.
-With a signal constructed, it remains to compute t-SNE outputs at varying perplexities.
-It’s recommended that at least a couple outputs are computed for each perplexity to account for t-SNE’s inherit randomness.
-For each output, one must calculate the trustworthiness and Shepard goodness with respect to the signal.
-From there, one can choose the representation with the desirable balance of local and global performance.
-A summary is given in Algorithm 1.
-Sample code is available at https://github.com/JustinMLin/DR-Framework/.
-Optimal Perplexity
-Dataset $n$ $p$ $r$ signal + noise signal
-Links [20] 500 10 3 30 110
-Trefoil [20] 500 10 3 30 110
-Mammoth [19] 500 10 3 20 100
-High-Dimensional Clusters 210 60 10 40 50
-CyTOF [15] 5,000 30 5 50 220
-CyTOF [15] 5,000 30 8 45 65
-scRNA-seq [16] 864 500 5 40 120
-scRNA-seq [16] 864 500 10 50 60
-Microbiome [2] 280 66 5 50 90
-Microbiome [2] 280 66 8 60 85
-Table 1: Summary of results
-It is worth noting that computational barriers may arise, especially for very large data sets.
-To alleviate such issues, trustworthiness and Shepard goodness can be approximated by subsampling before calculation.
-Furthermore, t-SNE is generally robust to small changes in perplexity [17], so checking a handful of different perplexities is sufficient.
-If computing the t-SNE representations is the limiting factor, the perplexity can be calibrated for a subsample of the data, instead.
-[14] found that embedding a $\rho$-sample, where $\rho\in(0,1]$ is the sampling rate, with perplexity $\textrm{Perp}^{\prime}$ gives a visual impression of embedding the original data with perplexity
-$\textrm{{Perp}}=\frac{\textrm{Perp}^{\prime}}{\rho}.$
-With these concessions, applying this framework to calibrate perplexity is feasible for data sets of any size.
 Algorithm 1 Measuring Performance in the Presence of Noise
@@
 12:Choose output with desired balance of local and global performance
+With a signal constructed, it remains to compute t-SNE/UMAP outputs at varying perplexities/n_neighbors.
+It’s recommended that at least a couple outputs are computed for each perplexity/n_neighbors to account for randomness in the algorithms.
+For each output, one must calculate the trustworthiness and Shepard goodness with respect to the signal.
+From there, one can choose the representation with the desired balance of local and global performance.
+A summary is given in Algorithm 1.
+Sample code is available at https://github.com/JustinMLin/DR-Framework/.
+It is worth noting that computational barriers may arise, especially for very large data sets.
+To alleviate such issues, trustworthiness and Shepard goodness can be approximated by subsampling before calculation.
+Furthermore, t-SNE and UMAP are generally robust to small changes in perplexity and n_neighbors, so checking a handful of values is sufficient.
+If computing multiple low-dimensional representations is the limiting factor, one can try calibrating the hyperparameters for a subsample before extending to the full data set.
+[24] found that embedding a $\rho$-sample, where $\rho\in(0,1]$ is the sampling rate, with perplexity $\textrm{Perp}^{\prime}$ gives a visual impression of embedding the original data with perplexity $\textrm{{Perp}}=\frac{\textrm{Perp}^{\prime}}{\rho}$.
+With these concessions, applying this framework to calibrate hyperparameters should be feasible for data sets of any size.
+6 Case Study
+To demonstrate how one might apply this framework, we walk through a detailed case study on a modern scRNA-seq data set.
+6.1 Data
+Cryopreserved human peripheral blood mononuclear cells (PBMCs) from a healthy female donor aged 25 were obtained by 10x Genomics from AllCells.
+Granulocytes were removed by cell sorting, followed by nuclei isolation.
+Paired ATAC and Gene Expression libraries were generated from the isolated nuclei and sequenced.
+See [25] for details.
+Figure 14: Scree Plot for PBMC Data Set
+Figure 15: Trustworthiness vs. n_neighbors for UMAP (PBMC)
+6.2 Pre-Processing
+Pre-processing was completed using the $R$ package BPCells and the steps followed the provided tutorial [26] closely.
+Low quality cells (those that did not meet the required number of RNA reads, the required number of ATAC reads, or TSS Enrichment cutoffs) were filtered out before a matrix normalization was applied.
+The cleaned dataset contained 2,600 cells and 1,000 genes.
+The number of dimensions was then reduced to 500 using PCA, which retained 86% of the original variance.
+The processed data set to be analyzed contained 2,600 observations in 500 dimensions, $\mathbb{Z+\epsilon}\in\mathbb{R}^{2,600\times 500}$.
+6.3 Determining the Signal
+To determine the number of signal dimensions, a scree plot was drawn (Figure 14).
+The first eigenvalue was approximately 188 but was trimmed to fit the plot.
+Four dimensions, a relatively conservative estimate, were chosen to represent the signal, $Y\in\mathbb{R}^{2,600\times 4}$.
+Figure 16: Cell Types (PBMC)
+Figure 17: Plot of Dendritic Cells (PBMC)
+6.4 Results
+UMAP was applied with multiple values of n_neighbors.
+20 representations were computed for each value, and trustworthiness was measured with respect to both the entire data and the signal.
+Trustworthiness was maximized at a n_neighbors value of 50 when comparing against the entire data and a value of 70 when comparing against the signal (Figure 15).
+Cell types (B, T, Monocyte, NK, Dendritic cell, CD8 T) were assigned to each cluster by exploring marker genes (Figure 16).
+See [26] for details.
+Figure 18: PCA Applied to Dendritic Cells (PBMC)
+Figure 19: PCA applied to Dendritic Cells (PBMC)
+6.5 Analysis
+In all three representations, the primary division of cell types is between monocytes and some of the dendritic cells vs.
+the T, B, CD8 T, and NK cells.
+In the default UMAP representation, the B cells form a cluster that is quite distinct from the T, CD8 T, and NK cells.
+As we increase n_neighbors to 50 and 70, the B cell cluster moves closer to the T/CD8 T/NK cell cluster.
+The closer proximity of the B cells to the T, CD8 T, and NK cells in the n_neighbors = 70 representation is consistent with the over-arching categorization of T, CD8 T, NK, and B cells as lymphocytes, as opposed to monocytes.
+Perhaps a starker difference between the representations concerns the dendritic cells (DCs).
+In the n_neighbors = 15 and n_neighbors = 50 representations, there are three distinct clusters of DCs, whereas there are only two in the n_neighbors = 70 representation (Figure 17).
+Principal component analysis of the DCs alone suggests that the DCs are either two clusters, one of which is more diffuse than the other, or three clusters, two of which are fairly close together (Figure 18).
+Standard metrics for determining the number of clusters suggest the same.
+The silhouette width metric suggests two clusters, while the gap statistic suggests three (Supporting Information SII.3).
+However, the three-cluster solution given by k-means and visual inspection of the principal components plot does not align with the three clusters in the n_neighbors = 15 or n_neighbors = 50 representation.
+The green and orange clusters are represented faithfully, but the third, more diffuse, purple cluster is split across two DC clusters in the n_neighbors = 15 and n_neighbors = 50 representations (Figure 19).
+The degree of separation is much lesser in the n_neighbors = 70 representation.
+Therefore, the n_neighbors = 15 and n_neighbors = 50 representations inaccurately represent the dendritic cells in a way that the n_neighbors = 70 representation does not.
 7 Discussion
@@
 When using the original data to calibrate perplexity, our experiments agreed with perplexities previously recommended.
-When using just the signal, however, our experiments indicated that larger perplexities performed better.
-Low perplexities lead to overly-flexible t-SNE models that are heavily impacted by the presence of noise, while higher perplexities exhibit better performance due to increased stability.
+When using the signal, however, our experiments indicated that larger perplexities perform better.
+Low perplexities/n_neighbors lead to overly-flexible models that are heavily impacted by the presence of noise, while higher perplexities/n_neighbors exhibit better performance due to increased stability.
 These considerations are especially important when working with heavily noised data, which are especially prevalent in the world of single-cell transcriptomics [27].
@@
 In the case when a specific signal structure is desired, this framework can be used to determine which dimension reduction method best preserves the desired structure.
-Further works should explore alternative methods for extracting the signal in way that preserves the desired structure.
+Further works should explore alternative methods for extracting the signal a in way that preserves the desired structure.
 8 Data Availability
 All data and code are freely available at https://github.com/JustinMLin/DR-Framework/.
-Supplementary Information
+Supporting Information
 SI Simulated Examples
 SI.1 Trefoil Plots
-Figure S1: Trefoil Plots
+For the trefoil example, the signal $Y$ consisted of a trefoil knot embedded in three dimensions containing 500 points.
+$Z+\epsilon$ was constructed by adding seven superfluous dimensions and isotropic Gaussian noise.
+Various degrees of noise were tested ($sd=5,10,15,20,25,30$).
+The first two plots depict Trustworthiness vs.
+Perplexity and the trustworthiness-maximizing embeddings for the $sd=10$ case.
+The third plot shows the trustworthiness-maximizing perplexity for the different degrees of noise.
+Figure S1: Trustworthiness vs. Perplexity (Trefoil)
+Figure S2: Trustworthiness-Maximizing Representations (Trefoil)
+Figure S3: Optimal Perplexity (Trefoil)
 SI.2 Mammoth Plots
-Figure S2: Mammoth Plots
+For the mammoth example, the signal $Y$ consisted of 500 points in three dimensions.
+The data was randomly sampled from the mammoth data set used in [18].
+$Z+\epsilon$ was constructed by adding seven superfluous dimensions and isotropic Gaussian noise.
+Various degrees of noise were tested ($sd=0.5,1,1.5,2,2.5,3$).
+The first two plots depict Trustworthiness vs.
+Perplexity and the trustworthiness-maximizing embeddings for the $sd=1$ case.
+The third plot shows the trustworthiness-maximizing perplexity for the different degrees of noise.
+Figure S4: Trustworthiness vs. Perplexity (Mammoth)
+Figure S5: Trustworthiness-Maximizing Representations (Mammoth)
+Figure S6: Optimal Perplexity (Mammoth)
 SII Practical Examples
-SII.1 UMAP Plots (CyTOF)
-Figure S3: UMAP Plots
-SII.2 scRNA Dataset
-This is a dataset of induced pluripotent stem cells generated from three different individuals [16].
-The original data includes 864 units and 19,027 readings per unit.
-To process this zero-inflated count data, columns containing a large proportion of 0’s (20% or more) were removed before a log transformation was applied.
-This reduced the dimension to 5,431.
-A PCA pre-processing stop further reduced the dimension to 500, which still retained 88% of the variance of the log-transformed data.
-The signal was first taken to be the first five principal components, then the first 10 principal components.
-Notice the optimal perplexity when compared against the original data differed between these two experiments, even though it should theoretically be independent of the chosen signal dimension.
-This is due to the inherent randomness of the t-SNE algorithm.
-Figure S4: scRNA Plots (r = 5)
-Figure S5: scRNA Plots (r = 10)
-SII.3 Microbiome Dataset
+SII.1 CyTOF Data Set
+The CyTOF data set contained 239,933 observations in 49 dimensions [21].
+To reduce the computational load, a subset of 5,000 observations was sampled.
+A log transformation was followed by a PCA pre-processing step to reduce the number of dimensions to 30, which still retained 77% of the variance in the original data.
+The processed data set to be studied consisted of 5,000 observations in 30 dimensions.
+The signal was first taken to be the first five principal components, then the first eight principal components.
+A hierarchical clustering of the high-dimensional data was computed then projected onto the trustworthiness-maximizing representations.
+Figure S7: Trustworthiness vs. Perplexity for r = 5 (CyTOF)
+Figure S8: Trustworthiness-Maximizing Representations for r = 5 (CyTOF)
+Figure S9: Trustworthiness vs. Perplexity for r = 8 (CyTOF)
+Figure S10: Trustworthiness-Maximizing Representations for r = 8 (CyTOF)
+SII.2 Microbiome Data Set
 [22] compares the faecal microbial communities from 22 subjects using complete shotgun DNA sequencing.
@@
 The signal was first taken to be the first five principal components, then the first eight principal components.
-Notice the optimal perplexity when compared against the original data differed between these two experiments, even though it should theoretically be independent of the chosen signal dimension.
-This is due to the inherent randomness of the t-SNE algorithm.
-Figure S6: Microbiome Plots (r = 5)
-Figure S7: Microbiome Plots (r = 8)
+A k-means clustering of the high-dimensional data was computed then projected onto the trustworthiness-maximizing representations.
+Figure S11: Trustworthiness vs. Perplexity for r = 5 (Microbiome)
+Figure S12: Trustworthiness-Maximizing Representations for r = 5 (Microbiome)
+Figure S13: Trustworthiness vs. Perplexity for r = 8 (Microbiome)
+Figure S14: Trustworthiness-Maximizing Representations for r = 8 (Microbiome)
+SIII PBMC Data Set
+Figure S15: Average Silhouette Width for Dendritic Cells
+Figure S16: Gap Statistic for Dendritic Cells